OpenStudy (osanseviero):

Write the next series completly and write to what tends when you get the limit

OpenStudy (osanseviero):

\[Sn=\frac{ 6 }{ 10 }+\frac{ 6 }{ 100 }+...\]

OpenStudy (anonymous):

is your job to add up \[\frac{6}{10}+\frac{6}{100}+\frac{6}{1000}+...\] which is the same as \[0.6666...\]

OpenStudy (anonymous):

you probably already know this one since \(0.33333...=\frac{1}{3}\) then if you double it you get \(0.6666...=\frac{2}{3}\)

OpenStudy (osanseviero):

Yep, I know that...so what should I do...keep adding?

OpenStudy (osanseviero):

oh....I see...it is 0.6666

OpenStudy (osanseviero):

So it tends to 0.66666... in infinity?

OpenStudy (osanseviero):

And how to write the complete series?

OpenStudy (anonymous):

\(\overline{.6}=\frac{2}{3}\)

OpenStudy (anonymous):

i was assuming you knew what "point six" repeating is if you have so sum a geometric series, we can do that as well, but you are still going to get \(\frac{2}{3}\)

OpenStudy (osanseviero):

I knew. So...its tendency and writing the series is the same?

OpenStudy (osanseviero):

Oh... \[Sn?\frac{ 6 }{ 10 }\times \frac{ 1 }{ 10 }^{n-1}\]

OpenStudy (osanseviero):

that is an =

OpenStudy (anonymous):

are you familiar with sigma notation?

OpenStudy (osanseviero):

Yep

OpenStudy (osanseviero):

so...